OpenStudy (anonymous):

d/dx [(e^2x-5)(secx)]???

OpenStudy (anonymous):

start with \[(e^{2x}-5)\cos(x)\] and use the product rule

OpenStudy (anonymous):

\[f'(x)=2e^{2x}\cos(x)-(e^{2x}-5)\sin(x)\]

OpenStudy (anonymous):

it's e^(2x-5)

OpenStudy (anonymous):

is the answer e^(2x-5)*secx*tanx+2*e^(2x-5)*secx